OpenStudy (anonymous):

Use the commutative law of addition to write an equivalent expression. ab + 3

OpenStudy (anonymous):

ab+3 = 3 + ab

OpenStudy (anonymous):

im not sure of ur answer

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

commuting is like getting on a bus. You can go wherever you want. addition and multiplication is does not matter what order you do them in. ab = ba x + y = y + x a + b + c + d = c + a + d + b so the smart aleky answer you your first question would be 3 + ba

OpenStudy (anonymous):

that makes sense